OpenStudy (anonymous):

what is the focus of x^2=20y

OpenStudy (anonymous):

compare it with the standard form x^2=4ay

OpenStudy (jack1):

x^2 = 20y x^2 = 4by so b = ...? @eric09 ? Vertex = O = ( 0, 0 ), Focus = ( 0, b )